A conical tent is 10 m high and the radius of its base is
24 m. Find
(i) slant height of the tent.
(u') cost of the canvas required to make the tent, if the cost of 1 m2 canvas is r 70.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Base radius = 24 m, height = 10 m. (i) Slant height = (24)2 + (10)2 m = 576 + 100 m = ./676 m = 26 m. (ii) 'lbtal canvas required = wl = x 24 >< 26 m2 Cost = (70x x 24 x 26) = 'f 137280.